Solve the following system of equations using any algebraic method
y=-3x+5
2x+y=6

Answer by macston(5194)

y=-3x+5
2x+y=6 Substitute for y
2x+(-3x+5)=6
-x+5=6 Subtract 5 from each side.
-x=1 Divide by -1
x=-1 ANSWER 1: x=-1
Using this, back solve equation 1 for y.
y=-3x+5
y=-3(-1)+5
y=3+5=8 ANSWER 2: y=8
CHECK:
Use second equation:
2x+y=6
2(-1)+8=6
-2+8=6
6=6